OpenStudy (anonymous):

solutions to 5x^2-2x+16=4x^2+6x

OpenStudy (anonymous):

is it -4?

OpenStudy (anonymous):

x=4,4 do u need full solution

OpenStudy (anonymous):

I don't think i do. just numbers. so 4 and -4? but isn't this a perfect square since root 16 is 4?

OpenStudy (anonymous):

5x^2-4x^2-2x-8x+16=0 x^2-8x+16=0 (x-4)(x-4)=0 therefore x=+4,+4

OpenStudy (anonymous):

so just 4 then?

OpenStudy (anonymous):

s just 4

OpenStudy (anonymous):

Here is the detailed solution:- \[5x ^{2}-4x ^{2}-2x-6x+16=0\] \[x ^{2}-8x+16=0\] \[x ^{2}-4x-4x+16=0\] Or you can get roots of any equation by this formula.. \[-b \pm \sqrt{b ^{2}-4ac}/2a when a \neq0\] you can x=-4,4

OpenStudy (anonymous):

Oops sorry +4 only